Brain block, can't remember how to do probability question

This forum made possible through the generous support of SDN members, donors, and sponsors. Thank you.
Get help with your application

Use all the free resources available to you from SDN: articles, guides, expert advising, forums discussions, and school research.

Adventure-time

Full Member
5+ Year Member
Advertisement - Members don't see this ad
Hello, I have a quick question regarding probabilities.

Lets say I'm about to shoot a firework at my chest.

The outside jacket is 93% effective in preventing burns. I also have an inner jacket that is 89% effective in preventing burns. Then finally, I have an inner shirt that is 70% effective in preventing burns.

With all these precautions, what would be the probability I would get burned?
 
I believe this is an AND question since you can get burned by all 3 jackets. Not sure if that’s correct but then it will be:
Probability of getting burned by outer jacket X probability of getting burned by inner jacket X probability of getting burned by shirt
= 0.07 X 0.11 X 0.3 = 0.00231. Is that the answer?
 
I believe this is an AND question since you can get burned by all 3 jackets. Not sure if that’s correct but then it will be:
Probability of getting burned by outer jacket X probability of getting burned by inner jacket X probability of getting burned by shirt
= 0.07 X 0.11 X 0.3 = 0.00231. Is that the answer?

I'm not sure if its an "And" question, but I do know that they want me to calculate the probability of being burned in with the individual probabilities of each jacket and shirt taken into account.

Ok, so I thought about it and this is what I did (which was the same as you, so I think we are correct?).

100-93 = 7.0/100 = 0.07
100-89 = 11/100 = 0.11
100-70 = 30/100 = 0.30

.07*.11*.30 = 0.00231 * 100 = 0.231%.

I think that we agree?